OpenStudy (anonymous):

how do I simplify (9/4)^-1/2

OpenStudy (anonymous):

(9/4)^-1/2 = (4/9)^1/2 = (2/3) .....I just took the square root of 4 and 9 which is 2 and 3 respectively Hope you understood

OpenStudy (anonymous):

that's correct according to mah calculator

OpenStudy (anonymous):

but you flipped them!

OpenStudy (anonymous):

so how would I do like, (-1/64)^-4/3?